# Steering Generative Models with Experimental Data for Protein Fitness Optimization

### v1

@@ Abstract
Protein fitness optimization involves finding a protein sequence that maximizes desired quantitative properties in a combinatorially large design space of possible sequences. Recent developments in steering protein generative models (e.g diffusion models, language models) offer a promising approach. However, by and large, past studies have optimized surrogate rewards and/or utilized large amounts of labeled data for steering, making it unclear how well existing methods perform and compare to each other in real-world optimization campaigns where fitness is measured by low-throughput wet-lab assays. In this study, we explore fitness optimization using small amounts (hundreds) of labeled sequence-fitness pairs and comprehensively evaluate strategies such as classifier guidance and posterior sampling for guiding generation from different discrete diffusion models of protein sequences. We also demonstrate how guidance can be integrated into adaptive sequence selection akin to Thompson sampling in Bayesian optimization, showing that plug-and-play guidance strategies offer advantages compared to alternatives such as reinforcement learning with protein language models.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Proteins, sequences of amino acids, can be optimized for useful properties such as binding affinity, catalytic activity, or stability, numerically quantified as “fitness.” However, the protein optimization problem is challenging: The design space of proteins is vast, as a protein of length $M$ can be constructed in $20^{M}$ different ways; only a negligible fraction of possible protein sequences are functional (Romero & Arnold, 2009); and most experimental wet-lab assays can only provide $10^{2}-10^{3}$ fitness labels per round. As a result, researchers often rely on directed evolution, an iterative process that incrementally improves protein fitness (Packer & Liu, 2015). In each round of directed evolution, a protein is mutated, the fitness of the variants is measured, and the most beneficial variant is retained for the next iteration. However, this approach can be slow (often only one mutation is accumulated in each round), and it is often inefficient–as it is limited to a local search of very similar protein sequences.
@@
Figure 1: Comparison of steered generation for protein optimization (SGPO) to other ML-assisted workflows for protein engineering. (A) SGPO involves initializing a generative prior model to sample sequences with high natural likelihoods and steering that model with assay-labeled fitness data. Optimization is difficult because the design space is massive, and the throughput of wet-lab fitness assays (Erlenmeyer flask icon) is low, so adaptive learning across multiple iterations is beneficial. Previous methods have utilized generative models such as (B) fully zero-shot methods that sample highly natural sequences but do not utilize labeled fitness data or (C) those that only utilize labeled fitness. (D) Alternatively, supervised approaches involve enumerating to calculate fitness predictions for all variants in a design space, limiting them to optimizing few residues (i.e. $N<9$).
@@
In recent years, there has been strong interest in developing machine learning (ML)-assisted methods to more efficiently optimize protein fitness (Yang et al., 2019; Wittmann et al., 2021a; Hie & Yang, 2022; Yang et al., 2024b). Many recent studies have focused on generative approaches that combine both unlabeled and labeled data for protein design. Broadly, these methods achieve conditional generation by steering generative priors of natural protein sequences (Freschlin et al., 2022) with fitness data, which can be incorporated into an adaptive optimization cycle (Hie & Yang, 2022). These Steered Generation for Protein Optimization (SGPO) methods are advantageous because they do not face the individual limitations of other methods (Fig. 1A, Table 1). Firstly, SGPO utilizes labeled data, which is important for fitness goals that deviate from natural function (e.g. enzymes being engineered for non-native activities (Arnold, 2018; Yang et al., 2025)), unlike previous studies utilizing only generative priors of natural sequences, e.g. for antibody design (Generative: Zero-Shot, Fig. 1B, Hie et al. 2023). Secondly, generative priors (Wu et al., 2021; Hsu et al., 2024) sample sequences with high evolutionary likelihoods and potentially higher fitness, which is advantageous compared to approaches only using labeled data (Brookes et al., 2019; Brookes & Listgarten, 2020; Song & Li, 2024; Stanton et al., 2022; Gupta & Zou, 2019; Hie & Yang, 2022) (Generative: Adaptive, Fig. 1C). Finally, SGPO is not limited to a small design space, unlike many supervised ML-assisted directed evolution (MLDE) approaches, which require enumerating to calculate fitness predictions for all variants in the design space (Supervised, Fig. 1D) (Wu et al., 2019; Wittmann et al., 2021b; Yang et al., 2025; Li et al., 2024a; Vornholt et al., 2024; Jiang et al., 2024; Hsu et al., 2022; Ding et al., 2024; Hawkins-Hooker et al., 2024; Zhao et al., 2024a; Thomas et al., 2025).
@@
Table 1: SGPO is a general approach for protein fitness optimization that doesn’t face the individual limitations of other strategies. Namely, SGPO utilizes zero-shot knowledge from the natural distribution of proteins, can be guided by assay-labeled fitness data, and can optimize many residues ($N$) simultaneously. Beyond those listed here, there are many other studies that combine different elements of these approaches.
@@
SGPO Lisanza et al. (2024); Widatalla et al. (2024); Stocco et al. (2024); Nisonoff et al. (2024); Blalock et al. (2025); Goel et al. (2024)
Generative: Zero-Shot $\times$ Hie et al. (2023)
Generative: Adaptive $\times$ Brookes et al. (2019); Brookes & Listgarten (2020); Song & Li (2024); Stanton et al. (2022)
Supervised $\times$ Wittmann et al. (2021b); Ding et al. (2024); Hawkins-Hooker et al. (2024); Zhao et al. (2024a)
Despite these advantages, SGPO methods remain limited in their applicability to real-world fitness optimization–specifically considering two classes of methods: guiding discrete diffusion models (Nisonoff et al., 2024; Stark et al., 2024; Klarner et al., 2024; Gruver et al., 2023; Lisanza et al., 2024; Goel et al., 2024) and reinforcement learning (RL) by finetuning models such as protein language models (PLMs) (Ruffolo & Madani, 2024; Widatalla et al., 2024; Stocco et al., 2024; Blalock et al., 2025). The limitations of prior work are summarized as follows: (1) Few previous studies have explored steering with few ($10^{2}-10^{3}$) labeled sequences (Lisanza et al., 2024; Stocco et al., 2024) for protein optimization based on real fitness data, e.g. activity or fluorescence, rather than computational surrogates (Lisanza et al., 2024; Blalock et al., 2025). (2) Most studies only evaluate one type of generative prior and steering strategy, so it is unclear how different combinations perform in practice. (3) Most existing approaches fail to incorporate principles from adaptive optimization, such as uncertainty-aware exploration (e.g., Bayesian optimization), which have shown clear benefits in protein engineering (Vornholt et al., 2024; Yang et al., 2025).
In this study, we focus on understanding the best practices for integrating SGPO into real-world wet-lab workflows. We focus on the following questions: Which steering strategies perform best, and with which types of models? How can we utilize uncertainty to better explore the design space when performing guidance? Overall, we make the following key contributions:
@@
    We comprehensively evaluate design decisions for SGPO, including different generative models for sequences and steering strategies (Fig. 3, Section 2), offering best practices for protein optimization with few fitness labels.
@@
    We adapt decoupled annealing posterior sampling (Zhang et al., 2024) and are the first to demonstrate the strong performance of such plug-and-play sampling techniques for SGPO.
On the TrpB and CreiLOV protein fitness datasets, we find that SGPO methods can consistently identify high-fitness protein variants. In particular, our results highlight the advantages of plug-and-play guidance with diffusion models over finetuned language models—offering greater steerability and lower computational cost. To support future research and real-world adoption, our extensive, user-friendly code is available at https://github.com/jsunn-y/SGPO.
@@ 2 Related work
Generative models for discrete sequences.
The most widely adopted generative models for natural protein sequences are PLMs, such as autoregressive transformers (Nijkamp et al., 2023) and masked language models (Rives et al., 2021). Increasingly, various diffusion model (Ho et al., 2020) architectures have shown efficacy for modeling discrete data (${\mathbf{x}}$), such as protein sequences (Alamdari et al., 2023; Wang et al., 2024b), leveraging many similar learning techniques such as masking or autoregressive decoding (Sahoo et al., 2024; Lou et al., 2024; Nie et al., 2025) (Fig. 2). These generative prior models $p({\mathbf{x}})$ can broadly be categorized into two types: those that perform diffusion in a continuous latent space (Li et al., 2022; Chen et al., 2023a; Dieleman et al., 2022) and those that diffuse directly over discrete space (Fig. 2). In the protein domain, it has also been shown that latent diffusion over embeddings from PLMs can be more effective (Meshchaninov et al., 2025; Chen et al., 2024; Torres et al., 2025). Alternatively, models performing diffusion in discrete space use a transition matrix to update all discrete states in each timestep (D3PM) (Austin et al., 2023), which has later been formulated as continuous-time Markov chains (Lou et al., 2024; Campbell et al., 2022, 2024; Schiff et al., 2024). There are two common ways to add noise to discrete sequences, using uniform noise matrices or absorbing state (masking) matrices (Fig. 2). These have been followed by simplified frameworks showing some of the highest performance for modeling natural language, such as masked diffusion language models (MDLMs) (Sahoo et al., 2024; Hoogeboom et al., 2022) and a variation that uses uniform noise called uniform diffusion language models (UDLMs) (Schiff et al., 2024). We elaborate more on these methods in Section A.3.
Figure 3: Methods design space for SGPO: a non-exhaustive landscape of generative models for protein sequences and methods to steer them with experimental labeled data. Three major types of diffusion models for sequences include those performing diffusion over continuous space and those performing diffusion over discrete space with a uniform or absorbing state (masking) noising process. Various types of guidance strategies are compatible with certain models, in green (NOS: diffusion optimization sampling, SMC: sequential monte carlo, FUDGE: future discriminators for generation, PPLM: plug and play language models, DDPP: discrete denoising posterior prediction, RTB: relative trajectory balance, DPLM: diffusion protein language model). Differently, language models and variational autoencoders can be aligned with labeled data via reinforcement learning such as policy optimization or supervised finetuning.
@@
An advantage of diffusion models is the ability to perform plug-and-play guidance based on fitness labels (${\mathbf{y}}$) without finetuning the generative prior model weights, resulting in reduced training costs and potentially strong signal despite having few ($\sim 10^{2})$ labels. Guiding a continuous diffusion model often involves skewing the learned score function using gradients from a supervised value function that can predict labels ${\mathbf{y}}$ from data ${\mathbf{x}}$ (Chung et al., 2023; Zheng et al., 2025; Soares et al., 2025). These methods are often referred to as posterior sampling, as they aim to sample from the posterior distribution, $p({\mathbf{x}}|{\mathbf{y}})$. Recent works extend this idea to guiding discrete diffusion models. Classifier guidance (CG) (Nisonoff et al., 2024) skews the rate matrix of the reverse time Markov chain of discrete diffusion models using a time-dependent value function, $p({\mathbf{y}}|{\mathbf{x}}_{t},t)$; variable splitting methods (DAPS) (Zhang et al., 2024; Chu et al., 2025) use discrete diffusion models as denoisers and only require a value function of clean data, $p({\mathbf{y}}|{\mathbf{x}}_{0})$; diffusion optimized sampling (NOS) (Gruver et al., 2023) trains a value function on continuous embeddings of discrete tokens and optimizes the embedding for higher fitness; sequential Monte Carlo methods (SMC) (Li et al., 2024b; Uehara et al., 2025; Wu et al., 2024; Lee et al., 2025a; Singhal et al., 2025) evolve multiple particles from a series of distributions to approximate the posterior distribution in limit. We explain these methods in more detail in Section A.4, along with other variations on the guidance process. In this study, we focus on CG, DAPS, and NOS as guidance techniques (Fig. 3). Future work could also consider guidance techniques for autoregressive language models, such as future discriminators for generation (FUDGE) (Yang & Klein, 2021), plug and play language models (PPLM) (Dathathri et al., 2020), and twisted SMC (Zhao et al., 2024b; Amin et al., 2024). Additionally, Xiong et al. (2025) demonstrate how guidance generalizes to masked language models and order-agnostic autoregressive models.
@@
We consider RL broadly here as enabling conditional generation by finetuning generative models with labeled data, thus pushing those models to produce more favorable generations. There are emerging RL techniques applied to discrete diffusion models, including discrete denoising posterior prediction (DDPP) (Rector-Brooks et al., 2024), relative trajectory balance (RTB) (Venkatraman et al., 2025), and direct reward backpropagation with gumbel softmax trick (DRAKES) (Wang et al., 2024a). While the above strategies are specific to discrete diffusion models, supervised fine-tuning (SFT) and policy optimization are two important techniques used in RL that can be broadly applied to generative models such as language models (Fig. 3). Policy optimization has generally shown better performance than SFT (Stocco et al., 2024; Blalock et al., 2025); in particular, direct preference optimization (DPO) is often used for its algorithmic simplicity and ease of training (Rafailov et al., 2023) (details in Section A.4). RL has demonstrated utility for aligning generative models of proteins (language models, inverse folding models, variational autoencoders) with properties like stability (Widatalla et al., 2024; Blalock et al., 2025; Stocco et al., 2024; Lim et al., 2025), but these methods can have high computational costs of finetuning and may require large amounts of labels ($>10^{3}$) to effectively steer generations. We include DPO with an autoregressive PLM (finetuned ProGen2 (Nijkamp et al., 2023)) as a baseline.
@@
Protein engineering is commonly conducted through adaptive workflows such as directed evolution Packer & Liu (2015) or ML-based approaches such as Bayesian optimization (Frazier, 2018; Stanton et al., 2022). These methods follow an iterative loop: labeled data is collected via expensive wet-lab assays, a surrogate model $p({\mathbf{y}}|{\mathbf{x}})$ is trained or updated, an acquisition function implied by the surrogate is used to propose new sequences to evaluate, and the cycle repeats (Hie & Yang, 2022; Vornholt et al., 2024; Yang et al., 2025). The surrogate model, often a Gaussian process or a deep ensemble, provides uncertainty estimates, which are used by an acquisition function (e.g. expected improvement, Thompson sampling) to balance exploration and exploitation of the design space. In this study, we adapt these ideas to guide diffusion models for protein sequence generation, as described in Section 4.3. A closely related line of work is latent space Bayesian optimization (Maus et al., 2022; Stanton et al., 2022; Lee et al., 2025b), which searches for optimal sequences within a latent space—typically learned by an autoencoder. However, it can be difficult to ensure that decoded sequences are valid and biologically meaningful (Maus et al., 2022).

### v2

@@ Abstract
Protein fitness optimization involves finding a protein sequence that maximizes desired quantitative properties in a combinatorially large design space of possible sequences. Recent advances in steering protein generative models (e.g., diffusion models and language models) with labeled data offer a promising approach. However, most previous studies have optimized surrogate rewards and/or utilized large amounts of labeled data for steering, making it unclear how well existing methods perform and compare to each other in real-world optimization campaigns where fitness is measured through low-throughput wet-lab assays. In this study, we explore fitness optimization using small amounts (hundreds) of labeled sequence-fitness pairs and comprehensively evaluate strategies such as classifier guidance and posterior sampling for guiding generation from different discrete diffusion models of protein sequences. We also demonstrate how guidance can be integrated into adaptive sequence selection akin to Thompson sampling in Bayesian optimization, showing that plug-and-play guidance strategies offer advantages over alternatives such as reinforcement learning with protein language models. Overall, we provide practical insights into how to effectively steer modern generative models for next-generation protein fitness optimization.
[…]
@@ 1 Introduction
Proteins, sequences of amino acids, can be optimized for useful properties such as binding affinity, catalytic activity, or stability, numerically quantified as “fitness.” However, protein optimization is challenging: the design space of proteins is enormous, as a protein of length $M$ can be constructed in $20^{M}$ different ways, of which only a negligible fraction are functional (Romero & Arnold, 2009). Moreover, most wet-lab assays only provide $10^{2}-10^{3}$ fitness labels. Consequently, researchers often rely on directed evolution, an iterative process aiming to incrementally improve protein fitness (Packer & Liu, 2015) through multiple rounds of mutation and experimental screening. In each round, a protein is mutated, the variants’ fitnesses are measured, and the most beneficial variant is selected for the next iteration. However, this approach can be slow, often accumulating only one mutation per round, and inefficient, as it performs a local search limited to closely related protein sequences.
@@
Figure 1: Comparison of steered generation for protein optimization (SGPO) to other ML-assisted workflows for protein engineering. (A) SGPO involves initializing a generative prior model to sample sequences with high natural likelihoods and steering that model with assay-labeled fitness data. Optimization is difficult because the design space is vast, and the throughput of wet-lab fitness assays (Erlenmeyer flask icon) is low, so adaptive learning across multiple iterations is beneficial. Previous methods have utilized generative models such as (B) fully zero-shot methods that sample highly natural sequences but do not utilize labeled fitness data or (C) those that only utilize labeled fitness. (D) Alternatively, supervised approaches involve enumerating to calculate fitness predictions for all variants in a design space, limiting them to optimizing few residues (i.e., $N<9$).
@@
In recent years, there has been a growing interest in developing machine learning (ML)-assisted methods to optimize protein fitness more efficiently (Yang et al., 2019; Wittmann et al., 2021a; Hie & Yang, 2022; Yang et al., 2024, 2025c). Many recent studies have focused on generative approaches combining unlabeled and labeled data for protein design. Broadly, these methods achieve conditional generation by steering generative priors of natural protein sequences (Freschlin et al., 2022) using fitness data, thereby enabling incorporation of the steered models into adaptive optimization cycles (Hie & Yang, 2022). We refer to this class of methods as Steered Generation for Protein Optimization (SGPO). These methods address the individual limitations of previous approaches (Fig. 1A, Table 1). First, SGPO leverages labeled data, which is essential for fitness goals that deviate from natural functions (e.g., engineering enzymes for non-native activities (Arnold, 2018; Yang et al., 2025b)), unlike zero-shot methods relying solely on generative priors of natural sequences (Generative: Zero-Shot, Fig. 1B, Hie et al. 2023; Sumida et al. 2024; Fei et al. 2025; Seki et al. 2025; Lambert et al. 2025). Second, generative priors (Wu et al., 2021; Hsu et al., 2024) sample sequences with high evolutionary likelihoods and potentially higher fitness, giving these methods a significant advantage over approaches relying exclusively on labeled data (Song & Li, 2023; Stanton et al., 2022; Gupta & Zou, 2019; Brookes & Listgarten, 2020; Jain et al., 2022; Kim et al., 2025; Angermueller et al., 2020; Hie & Yang, 2022) (Generative: Adaptive, Fig. 1C). Finally, SGPO scales to larger design spaces, unlike most supervised ML-assisted directed evolution (MLDE) approaches, which require enumerating and scoring all variants in the design space (Supervised, Fig. 1D) (Wu et al., 2019; Wittmann et al., 2021b; Yang et al., 2025b; Li et al., 2025a; Vornholt et al., 2024; Jiang et al., 2024; Hsu et al., 2022; Ding et al., 2024; Hawkins-Hooker et al., 2024; Zhao et al., 2024a; Thomas et al., 2025; Sun et al., 2025).
Despite these advantages, SGPO methods still face practical limitations in real-world fitness optimization, particularly across two major classes of approaches: guiding discrete diffusion models (Nisonoff et al., 2025; Stark et al., 2024; Klarner et al., 2024; Gruver et al., 2023; Lisanza et al., 2025; Goel et al., 2025) and finetuning models such as protein language models (PLMs) through reinforcement learning (RL) (Ruffolo & Madani, 2024; Widatalla et al., 2024; Stocco et al., 2024; Blalock et al., 2025; Wang et al., 2025c). The limitations of prior work are summarized as follows: (1) Few previous studies have explored steering with few ($10^{2}-10^{3}$) labeled sequences (Lisanza et al., 2025; Stocco et al., 2024) for protein optimization based on real fitness data, e.g. activity or fluorescence, rather than computational surrogates (Lisanza et al., 2025; Blalock et al., 2025). (2) Most studies only evaluate one type of generative prior and steering strategy, so it is unclear how different combinations perform in practice. (3) There is room to incorporate principles from adaptive optimization, such as uncertainty-aware exploration (e.g., Bayesian optimization), which have shown clear benefits in protein engineering (Vornholt et al., 2024; Yang et al., 2025b).
@@
Table 1: SGPO is a general approach for protein fitness optimization that does not face the individual limitations of other strategies. Namely, SGPO utilizes zero-shot knowledge from the natural distribution of proteins, can be guided by assay-labeled fitness data, and can optimize many residues (NN) simultaneously. Beyond those listed here, there are many other studies that combine different elements of these approaches.
@@
SGPO Lisanza et al. (2025); Widatalla et al. (2024); Stocco et al. (2024); Nisonoff et al. (2025); Brookes et al. (2019); Blalock et al. (2025); Goel et al. (2025); Huang et al. (2025)
Generative: Zero-Shot $\times$ Hie et al. (2023); Sumida et al. (2024); Fei et al. (2025); Seki et al. (2025); Lambert et al. (2025)
Generative: Adaptive $\times$ Song & Li (2023); Jain et al. (2022); Angermueller et al. (2020); Stanton et al. (2022); Brookes & Listgarten (2020)
Supervised $\times$ Wittmann et al. (2021b); Ding et al. (2024); Hawkins-Hooker et al. (2024); Zhao et al. (2024a); Sun et al. (2025)
In this study, we aim to understand the best practices for integrating SGPO into real-world engineering workflows. We focus here on modern generative models (i.e., discrete diffusion, language models) but acknowledge that other related methods are relevant, such as those based on variational autoencoders (Brookes et al., 2019; Torres et al., 2024) and other adaptive search strategies (Kirjner et al., 2024; Sinai et al., 2020; Ren et al., 2022). We explore the following questions: Which steering strategies perform best, and with which types of models? How can we utilize uncertainty to better explore the design space when performing guidance? Overall, we make the following key contributions:
@@
    We comprehensively evaluate design decisions for SGPO, including different generative models for sequences and steering strategies (Fig. 2 & 3, Section 2), offering best practices for protein optimization with few fitness labels.
@@
    We are the first to adapt decoupled annealing posterior sampling (Zhang et al., 2025) for SGPO, and this type of plug-and-play guidance has the strongest performance overall.
On the TrpB, CreiLOV, and GB1 protein fitness datasets, we find that SGPO methods can consistently identify high-fitness protein variants. In particular, our results highlight the advantages of plug-and-play guidance with diffusion models over finetuned language models—offering greater steerability and lower computational cost. To support future research and real-world adoption, our extensive, user-friendly code is available at https://github.com/jsunn-y/SGPO.
@@ 2 Related work
Generative models for discrete sequences.
The most widely adopted generative models for natural protein sequences are PLMs, such as autoregressive transformers (Nijkamp et al., 2023) and masked language models (Rives et al., 2021). Increasingly, various diffusion model (Ho et al., 2020) architectures have shown efficacy for modeling discrete data ($\mathbf{x}$}) (Li et al., 2025b), such as protein sequences (Alamdari et al., 2024; Wang et al., 2024), leveraging many similar learning techniques such as masking or autoregressive decoding (Sahoo et al., 2024; Lou et al., 2024; Nie et al., 2025; Shi et al., 2024) (Fig. 2). These generative prior models $p({\mathbf{x}})$ can be categorized broadly into two types: those that perform diffusion in a continuous latent space (Li et al., 2022; Chen et al., 2023b; Dieleman et al., 2022) and those that diffuse directly over discrete space (Fig. 2). In the protein domain, it has also been shown that latent diffusion over embeddings from PLMs can be more effective (Meshchaninov et al., 2025; Chen et al., 2023a; Torres et al., 2025). Alternatively, models performing diffusion in discrete space use a transition matrix to update all discrete states in each timestep (D3PM) (Austin et al., 2021), which has later been formulated as continuous-time Markov chains (Lou et al., 2024; Campbell et al., 2022, 2024; Schiff et al., 2024). Two common ways to add noise to discrete sequences are to use uniform noise matrices or absorbing state (masking) matrices (Fig. 2). These have been followed by simplified frameworks showing some of the highest performance for modeling natural language, such as masked diffusion language models (MDLMs) (Sahoo et al., 2024; Hoogeboom et al., 2022; Shi et al., 2024) and a variation that uses uniform noise called uniform diffusion language models (UDLMs) (Schiff et al., 2024). We elaborate more on these methods in Section A.3.
@@
An advantage of diffusion models is the ability to perform plug-and-play guidance based on fitness labels ($\mathbf{y}$}) without finetuning the generative prior model weights, resulting in reduced training costs and potentially strong signal despite having few ($\sim 10^{2})$ labels. Guiding a continuous diffusion model often involves skewing the learned score function using gradients from a supervised value function that can predict labels $\mathbf{y}$} from data $\mathbf{x}$} (Chung et al., 2023; Zheng et al., 2025; Soares et al., 2025). These methods are often referred to as posterior sampling, as they aim to sample from the posterior distribution, $p({\mathbf{x}}|{\mathbf{y}})$. Recent works extend this idea to guiding discrete diffusion models. Classifier guidance (CG) (Nisonoff et al., 2025) skews the rate matrix of the reverse time Markov chain of discrete diffusion models using a time-dependent value function, $p({\mathbf{y}}|{\mathbf{x}}_{t},t)$; variable splitting methods (DAPS) (Zhang et al., 2025; Chu et al., 2025) use discrete diffusion models as denoisers and only require a value function of clean data, $p({\mathbf{y}}|{\mathbf{x}}_{0})$; diffusion optimized sampling (NOS) (Gruver et al., 2023) trains a value function on continuous embeddings of discrete tokens and optimizes the embedding for higher fitness; sequential Monte Carlo methods (SMC) (Li et al., 2024a; Uehara et al., 2025; Wu et al., 2024; Lee et al., 2025a; Singhal et al., 2025) evolve multiple particles from a series of distributions to approximate the posterior distribution in limit. We explain these methods in more detail in Section A.4, along with other variations on the guidance process. In this study, we focus on CG, DAPS, and NOS as guidance techniques (Fig. 3). Future work could also consider guidance techniques for autoregressive language models, such as future discriminators for generation (FUDGE) (Yang & Klein, 2021), plug and play language models (PPLM) (Dathathri et al., 2020), and twisted SMC (Zhao et al., 2024b; Amin et al., 2025b). Additionally, Xiong et al. (2025) demonstrate how guidance generalizes to masked language models and order-agnostic autoregressive models.
Figure 3: Methods design space for SGPO: a non-exhaustive landscape of generative models for protein sequences and methods to steer them with labeled data. Three major types of diffusion models for sequences include those that perform diffusion over continuous space and those that perform diffusion over discrete space with a uniform or absorbing state (masking) noising process. Various types of guidance strategies are compatible with certain models, in green (NOS: diffusion optimization sampling, SMC: sequential monte carlo, FUDGE: future discriminators for generation, PPLM: plug and play language models, DDPP: discrete denoising posterior prediction, RTB: relative trajectory balance, DPLM: diffusion protein language model, BO: Bayesian optimization). Differently, language models and variational autoencoders can be aligned with labeled data via reinforcement learning such as policy optimization or supervised finetuning.
@@
We consider RL broadly here as techniques that achieve conditional generation by finetuning generative models with labeled data, thus pushing those models to produce more favorable generations. There are emerging RL techniques applied to discrete diffusion models, including discrete denoising posterior prediction (DDPP) (Rector-Brooks et al., 2025), relative trajectory balance (RTB) (Venkatraman et al., 2024; Bartoldson et al., 2025; Venkatraman et al., 2025), and direct reward backpropagation with gumbel softmax trick (DRAKES) (Wang et al., 2025a). While the above strategies are specific to discrete diffusion models, supervised fine-tuning (SFT) and policy optimization are two important techniques used in RL that can be broadly applied to generative models such as language models (Fig. 3). Policy optimization has generally shown better performance than SFT (Stocco et al., 2024; Blalock et al., 2025); in particular, direct preference optimization (DPO) is often used for its algorithmic simplicity and ease of training (Rafailov et al., 2023) (details in Section A.4). RL has demonstrated utility for aligning generative models of proteins (language models, inverse folding models, variational autoencoders) with properties like stability (Widatalla et al., 2024; Blalock et al., 2025; Stocco et al., 2024; Lim et al., 2025), but these methods can have high computational costs of finetuning and may require large amounts of labels ($10^{3}$) to effectively steer generations. We include DPO with an autoregressive PLM (finetuned ProGen2 (Nijkamp et al., 2023)) as a baseline.
@@
Protein engineering is commonly conducted through adaptive workflows such as directed evolution Packer & Liu (2015) or ML-based approaches such as Bayesian optimization (Frazier, 2018; Stanton et al., 2022). These methods follow an iterative loop: labeled data is collected via expensive wet-lab assays, a surrogate model $p({\mathbf{y}}|{\mathbf{x}})$ is trained or updated, an acquisition function implied by the surrogate is used to propose new sequences to evaluate, and the cycle repeats (Hie & Yang, 2022; Vornholt et al., 2024; Yang et al., 2025b). The surrogate model, often a Gaussian process or a deep ensemble, provides uncertainty estimates, which are used by an acquisition function (e.g., expected improvement, Thompson sampling) to balance exploration and exploitation of the design space. In this study, we adapt these ideas to guide diffusion models for protein sequence generation, as described in Section 4.3. A closely related line of work is latent space Bayesian optimization (Maus et al., 2022; Stanton et al., 2022; Gómez-Bombarelli et al., 2018; Castro et al., 2022; Torres et al., 2024; Lee et al., 2025b), which searches for optimal sequences within a latent space—typically learned by an autoencoder, which can implicitly capture a prior on natural protein sequences. In this work, we compare against APEXGo (Torres et al., 2024), a method that performs trust-region Bayesian optimization in the latent space of a variational autoencoder trained over protein sequences. There are also related methods that involve conditional sampling from a prior (Brookes et al., 2019). However, we note that SGPO offers greater flexibility by avoiding reliance on an explicit latent space, which enables the use of modern, more powerful generative models such as diffusion models and protein language models that are not easily accommodated by traditional latent Bayesian optimization pipelines.